# Rise of the forsaken relics: connecting present-day stellar streams and phase-mixed galaxies to the Epoch of Reionization

### v2

@@ VI.2.1 Lowering the low-mass cutoff at $z=7$
We select progenitor galaxies with $M_{\star}\geq 10^{4.5}\,M_{\odot}$ instead of $M_{\star}\geq 10^{5}\,M_{\odot}$) at $z=7$. This increases the number of progenitor galaxies, which naturally leads to a smaller standard deviation on $\alpha$ in each simulation, for each of the samples. In this case, sample III produces the lowest standard deviation followed by sample II and I, respectively. However, we get steeper mean $\alpha$ values in the all progenitors scenario and flatter values in the fossil record scenario. This leads to a larger difference between the slopes in these two scenarios when more disrupted galaxies are added, even though the halo-to-halo scatter decreases. Our estimates using bootstrapping, shown in Table 2, demonstrate these effects. With a lower mass cutoff, Sample III provides the most accurate and precise estimate of the proto-MW/LG slope in the all progenitors scenario followed by sample II and I, respectively. However, in the fossil record scenario, sample I provides a more accurate representation of the proto-MW/LG slope than sample II and III. Thus reducing the mass cut-off incorporates more low-mass progenitors which leads to the steepening of the slope in the all progenitors scenario. However, when combined into a single progenitor in the fossil record scenario, we get flatter SMFs at the low-mass end. The effect gets more drastic when we add more massive present-day substructures since we increase the amplitude more towards the high-mass-end than in the low-mass-end.
VI.2.2 Increasing the present-day stellar mass cutoff for intact galaxies
We can reliably select present-day intact galaxies down to $M_{\star}\gtrsim 10^{4.5}\,M_{\odot}$, but disrupted galaxies only down to $M_{\star}\gtrsim 10^{6}\,M_{\odot}$. So we test our results by also selecting intact present-day galaxies with $M_{\star}\gtrsim 10^{6}\,M_{\odot}$. When we use progenitor galaxies with $M_{\star}\gtrsim 10^{4.5}\,M_{\odot}$ at high redshifts, we find similar results for the proto-MW/LG slope in the three samples, with the least halo-to-halo scatter in case of sample III in the all progenitors scenario. However, in the fossil record scenario all the three samples give a flatter slope than that of the proto-MW/LG. Upon increasing the cut-off to $M_{\star}\gtrsim 10^{5}\,M_{\odot}$ at high redshifts, the difference between the slopes in the all progenitors scenario and the fossil record scenario reduces, as expected. In the all progenitors scenario the slope of sample III is closer to the proto-MW/LG slope ($\Delta\tilde{\mu}_{\alpha}\approx 0.006$) than sample I ($\Delta\tilde{\mu}_{\alpha}\approx 0.145$). However, in the fossil record scenario, sample I is more accurate ($\Delta\tilde{\mu}_{\alpha}\approx 0.007$) than sample III ($\Delta\tilde{\mu}_{\alpha}\approx 0.150$).
@@ VI.2.3 Removing morphologically complex and extremely massive present-day substructures
Our present-day sample includes a few morphologically-complex, massive galaxies with $M_{\star}\gtrsim 10^{9}\,M_{\odot}$. These typically have $\sim 30-50$ progenitors and lead to more bias in the fossil record scenario. Hence we test our results by restricting our present-day galaxy sample to those with stellar masses between $10^{6}\,M_{\odot}\lesssim M_{\star}\lesssim 10^{9}\,M_{\odot}$. The trends in this case resemble our previous results without a high-mass cut-off.
@@ VI.2.4 Lowering the high-mass cutoff at the present day
We tried lowering the high-mass cutoff for present-day substructures to $M_{\star}\lesssim 10^{8}\,M_{\odot}$ while keeping the low-mass cut-off at $M_{\star}\gtrsim 10^{4.5}\,M_{\odot}$. In this case, we find that the slopes across the three samples in the all progenitors scenario, although similar, are steeper than that of the proto-MW/LG by $\approx 0.23$. In the fossil record scenario, on the other hand, the slopes are flatter than the proto-MW/LG; with sample III being flatter (by $\approx$0.4) than sample II (by $\approx$0.3) or sample I (by $\approx$0.2). Interestingly, the halo-to-halo scatter reduces only by $\sim$3% in sample II and by $\sim$15% in sample III. This remains similar when selecting progenitor galaxies with $M_{\star}\gtrsim 10^{5}\,M_{\odot}$ at high redshifts, although in this case all three samples produce a steeper slope than the proto-MW/LG (by  $\approx\,0.1-0.3$; sample III is more accurate than sample I).
@@ VI.2.5 Discussion of sensitivity analysis
Except when decreasing the upper mass limit, these results show that adding disrupted galaxies robustly reduces the halo-to-halo scatter, with the quantitative benefit depending on the sample selections employed. However, we can get a biased estimate of the slope in the fossil record approach, while still reducing the halo-to-halo scatter by adding more disrupted galaxies. The bias exists even while using just the intact galaxies, with the precise bias depending, again, on sample selection. The all progenitors scenario provides more accurate and precise recoveries of the underlying proto-MW/LG slope when disrupted galaxies are added. This motivates future efforts to combine the traditional fossil record analysis with simulated merger trees, which may allow one to properly account for the multiple high redshift progenitors of individual present-day substructures, both intact and disrupted. If, instead, only the fossil record is used, it is important to test sensitivity to the selection criteria used for both present-day and high-$z$ galaxies.
@@ VI.3 Redshift evolution
Figure 11: Redshift evolution of the low-mass/faint-end slope of the SMF/UVLF. Hierarchical assembly is reflected in the increasingly steep low-mass/faint-end slopes with redshift for all three samples, as well as the proto-MW/LG (black) and FIREboxHR (yellow), for the differential SMF (top) and UVLF (bottom). In both panels, black, red, blue, and green points show $\tilde{\mu}_{\alpha}$ values for the proto-MW/LG, sample I, II, and III respectively; errorbars show the $\tilde{\sigma}_{\alpha}$ values as a measure of halo-to-halo scatter. Empty markers indicate slopes estimated in the all progenitors scenario, while filled markers denote the fossil record case.
@@
We also estimate $\alpha$ for the SMFs/UVLFs at $z=$ 6, 8, and 9, using our fiducial choice of $M_{\star}\geq 10^{5}\,M_{\odot}$ or $M_{UV}\leq-10$ at high redshifts. We show the $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$ statistics as our central estimates for the low-mass/faint-end slope and the halo-to-halo scatter, respectively. The top panel of Figure 11 shows results for the low-mass-end slopes of the SMFs and the bottom panel shows the faint-end slopes for the UVLFs. The slope estimates from the all progenitors scenario is shown by the empty markers and dashed errorbars, while the filled markers and solid errorbars denote the slope estimates from the fossil record scenario. As a consequence of hierarchical structure formation in the cold dark matter picture with standard radiative and supernova feedback, the SMFs/UVLFs should generally get steeper towards higher redshifts. This is well-reflected by the SMFs/UVLFs from FIREboxHR (yellow), the proto-MW/LG (black), and all the three samples (I - red; II - blue; III - green).
@@
In general, sample III consistently reproduces the most accurate and precise proto-MW/LG slope across the redshifts followed by sample II and I, in the all progenitors scenario. The trend is similar in the fossil record scenario, except in a few cases. The reduction in the halo-to-halo scatter with sample II and III is evident across the range of redshifts shown. In particular, at $z=$ 8 and 9, the large scatter in the slope from the progenitors of the intact galaxies alone is fairly well mitigated upon adding the disrupted galaxies, as shown by sample III. We caution, however, that the $z=8$ and $z=9$ results may be subject to the limited statistical samples of progenitor galaxies captured in our simulations at these redshifts. At $z=$ 9, the faint-end slopes from our reconstructed UVLFs are shallower than the universal one. These estimates are probably poor because of the low halo abundance at these redshifts. Overall, for both the SMFs/UVLFs and for most redshifts, sample III reproduces the proto-MW/LG slope more accurately and also reduces the halo-to-halo scatter, followed by sample II and I. This is true for both the all progenitors and fossil record scenarios using our fiducial selection criteria. However, as discussed in section VI.1, the results can vary depending on our sample selection criteria, both at the present-day and at high redshifts.
VII Reconstructed cosmic UV density
@@
For each of the three samples in each halo, we compute the reconstructed cosmic UV density and explore their contribution to the total cosmic UV density. As shown in Figure 12 (top panel), the proto-MW/LG (which also includes the progenitors of the present-day main host galaxy), can recover between $\sim$18% and $\sim$45% of the total universal cosmic UV density. We cannot fully recover the total universal cosmic UV density because of the difference in normalization between the UVLFs here and that of FIREboxHR, as previously discussed in section V. Hence, we mainly compare the contribution from the three samples and the proto-MW/LG. Sample III ($\sim$8–18%) is the most complete, followed by sample II ($\sim$4–16%), and sample I ($\sim$1–8%).
We also consider the relative contribution of different progenitor galaxies to the total reconstructed cosmic UV density within each sample at $z=7$ (Figure 12, bottom panel). The median UV luminosity density reconstructed from sample I (red) under-predicts the contributions from galaxies brighter than $M_{\mathrm{UV}}\sim-15$, which is the current JWST detection limit for unlensed galaxies  [^5] Measurements behind foreground lensing clusters can probe fainter galaxies down to $M_{\mathrm{UV}}\gtrsim-13$ (Chemerynska et al., 2025), but they come with systematic uncertainties.. Furthermore, the large scatter in the faint-end slope of the reconstructed UVLF from sample I also leads to sizable uncertainties in the inferred UV luminosity density and the relative importance of different galaxies in reionization. Quantitatively, as inferred from sample I, the luminosity above which galaxies produce $\sim 50\%$ of the total UV density is somewhere in the range $16\lesssim M_{\mathrm{UV}}\lesssim-13.5$. Adding the progenitors of the disrupted galaxies in samples II and III (blue and green shaded regions, respectively) reduces the scatter. Sample II predicts minimum UV magnitudes in the range $16.5\lesssim M_{\mathrm{UV}}\lesssim-14.5$, while sample III predicts a range of $16.5\lesssim M_{\mathrm{UV}}\lesssim-15$. Furthermore, the median curves (solid curves) are more representative of the proto-MW/LG curve (black) when the disrupted galaxies are included. The median curve for sample III (green) shows slightly larger contributions from brighter galaxies than in the proto-MW/LG, because it predicts a shallower faint-end slope than that of the proto-MW/LG in the fossil record scenario (see Figure 10).
@@ VIII Detectability of the present-day substructures
To complete our analysis, we construct UVLFs by imposing the detectability constraints obtained from the internal view scenario using LSST. Figure 4—first shown in Section V.1—gives the differential UVLFs at $z=7$ reconstructed from different classification groups (solid curves) in comparison to reconstructions from their detectable counterparts alone (dashed curves). The reconstructed UVLFs appear promising, even after imposing the detectability constraints. Although the normalization of the satellite and field galaxy progenitor UVLFs is reduced, their shapes are relatively well-preserved. The progenitor UVLFs reconstructed from the streams remain almost unchanged even with just $\sim$50 % detectability, since the detectable streams are mostly of higher mass and each massive stream typically has numerous reionization-era progenitors (see Section IV.2), spanning much of the high-$z$ distribution in progenitor luminosity. We show the results in the all progenitors scenario by utilizing the full merger histories of the present-day substructures, and not in the fossil record scenario. However, the difference between the shape of the UVLFs with and without detectability constraints does not change significantly in the fossil record scenario.
@@ IX Summary
In this work, we test whether including present-day disrupted galaxies like stellar streams and phase-mixed galaxies in the near-far technique improves the inference of the low-mass/faint-end slope of the SMF/UVLF during the EoR. Using the FIRE-2 cosmological zoom-in baryonic simulations of MW-like (isolated) and LG-like (paired) halos (Section II), we classify present-day substructures into 4 categories: intact satellite galaxies, field galaxies, stellar streams, and phase-mixed galaxies (Section III). The first two categories constitute our intact galaxies sample (sample I); the intact galaxies and streams comprise our sample II; adding the phase-mixed galaxies makes our sample III. Using particle tracking, we assign each star particle to its progenitor galaxy at high redshifts. We then assess the contribution of different progenitor populations to the total UV luminosity density (Sections IVVI). Finally, we estimate the proportion of substructures that will be detected by LSST (for the Milky Way) or Roman (for external galaxies) and the effect of these detection limits on the reconstructions (Section VIII). Our results are summarized below.
Our results are consistent in general with Gandhi et al. (2024), with a few differences. Gandhi et al. (2024) estimated the low-mass-end slopes of the reconstructed fossil record SMFs using just intact low-mass galaxies and found close agreement with the proto-MW/LG slopes. We reach similar conclusions but also show that the halo-to-halo scatter associated with the slope is large when using just intact galaxies. Their slope estimates are not exactly the same as our estimates for sample I, since some of their present-day low-mass galaxies are part of our streams and hence do not show up in our sample I. This also leads to differences in the shape and normalization of our respective intact low-mass galaxy fossil record SMFs, with a stronger mass dependent bias in our case. Additionally, our methods for computing the slopes are different. Furthermore, Gandhi et al. (in prep.), use an alternative set of zoom-in high-redshift FIRE-2 simulations as a benchmark for the universal SMFs/UVLFs, instead of FIREboxHR. Both sets of simulations show good agreement with each other, as well as with observed high-$z$ SMFs/UVLFs and cosmic UV luminosity density, but they may not match exactly.
@@
Inclusion of tidally disrupted galaxies leads to more accurate high-$z$ reconstructed SMFs/UVLFs. We analyze the SMFs/UVLFs under two scenarios: the all progenitors scenario, in which we account for the stellar mass contribution from each progenitor of a present-day substructure separately; and the fossil record scenario, where we assume that the stellar mass from all progenitors of each substructure is contained within a single galaxy, as would be inferred observationally in the absence of merger history information. In both of these scenarios, our sample III recovers most of the amplitude ($\sim$75%) of the proto-MW/LG SMF/UVLF, followed by sample II ($\sim$50%) and I ($\sim$25%), respectively. This represents a significant improvement, reducing the normalization rescaling from a factor of $\sim$4 in sample I to just $\sim$1.5 in sample III, after including the disrupted galaxies. However, in the fossil record scenario, the SMFs/UVLFs become shallower in the low-mass/faint-end region, with the effect being large when more disrupted galaxies are added. Nevertheless, in most instances, the SMFs and UVLFs in the fossil record scenario trace the shape of the proto-MW/LG SMFs/UVLFs between $10^{5}\,M_{\odot}\lesssim M_{\star}\lesssim 10^{6.5}\,M_{\odot}$ and $14\lesssim M_{\mathrm{UV}}\lesssim-10$, respectively. The fossil record results differ from the all progenitors case because the latter approach combines multiple low-mass/faint progenitors into a single more massive system. This difference is compounded in the case of disrupted galaxies which are generally more massive and have multiple reionization-progenitors. We also compare our results with the differential SMFs/UVLFs of the FIREboxHR simulation, as a benchmark for the universal SMF/UVLF, and find close agreement in shape. Although the proto-MW/LG slopes are in agreement with FIREboxHR within our estimated uncertainties, they are in general flatter by $\sim$0.07. This could be due the difference in the mean matter density of the region around the MW/LG and the mean matter density of the universe, but a detailed exploration is needed, which is beyond the scope of this paper.
@@
The low-mass/faint-end slope is more robustly constrained when including disrupted galaxies. We estimate the low-mass/faint-end slope of the reconstructed SMFs/UVLFs using a Bayesian approach, to test how accurately and precisely we can recover the proto-MW/LG and the universal SMF/UVLF slopes. In the all progenitors scenario, adding disrupted galaxies helps reproduce the low-mass/faint-end slope of the proto-MW/LG SMF/UVLF more accurately and more precisely, and these are also representative of the global-average low-mass/faint-end slopes. Sample III reproduces the proto-MW/LG slope most accurately with an offset of $\approx$0.04, followed by sample II ($\approx$0.12) and I ($\approx$0.26). Furthermore, when we use only the intact galaxies, there is a large halo-to-halo scatter due to the varying accretion histories across different halos, even though the median values agree well with that from the proto-MW/LG. Sample III reduces the scatter of sample I by $\sim$40%, while sample II reduces it by $\sim$20%. Including the disrupted galaxies improves our inferences of the median slopes and also reduces the halo-to-halo scatter. In the fossil record scenario, we find similar results when we select high-$z$ progenitors with $M_{\star}\geq 10^{5}\,M_{\odot}$ and $M_{\mathrm{UV}}\leq-10$. However, even though the halo-to-halo scatter decreases, we can get a biased estimate of the slope from the fossil record depending on our choice of both the present-day and high-$z$ galaxy samples. The estimated slopes vary depending on the stellar mass range of the galaxies selected, both at the present-day and during the EoR, even in the case of using only intact galaxies. In the future, it may be possible to augment the observed SFHs with average merger histories from cosmological simulations, providing yet more accurate and precise estimates of the low-mass/faint-end slope. Nevertheless, including the streams and phase-mixed galaxies improves the statistical constraints on the proto-MW/LG SMFs/UVLFs, reducing the halo-to-halo scatter ($\sim$0.13 in case of proto-MW/LG) from $\sim$0.24 (sample I) to $\sim$0.20 (sample II) and $\sim$0.15 (sample III). A similar trend holds when we separate into MW-like and LG-like environment. This improvement is therefore crucial, since we have only one MW and LG with which to apply the near-far technique.

### v3

@@ VI.2.1 Lowering the Low-mass Cutoff at $z=7$
We select progenitor galaxies with $M_{\star}\geq 10^{4.5}\,M_{\odot}$ instead of $M_{\star}\geq 10^{5}\,M_{\odot}$) at $z=7$. This increases the number of progenitor galaxies, which naturally leads to a smaller standard deviation on $\alpha$ in each simulation, for each of the samples. In this case, Sample III produces the lowest standard deviation followed by Sample II and I, respectively. However, we get steeper mean $\alpha$ values in the all-progenitors scenario and flatter values in the fossil-record scenario. This leads to a larger difference between the slopes in these two scenarios when more disrupted galaxies are added, even though the halo-to-halo scatter decreases. Our estimates using bootstrapping, shown in Table 2, demonstrate these effects. With a lower-mass cutoff, Sample III provides the most accurate and precise estimate of the proto-MW/LG slope in the all-progenitors scenario followed by Samples II and I, respectively. However, in the fossil-record scenario, Sample I provides a more accurate representation of the proto-MW/LG slope than Samples II and III. Thus reducing the mass cut-off incorporates more low-mass progenitors which leads to the steepening of the slope in the all-progenitors scenario. However, when combined into a single progenitor in the fossil-record scenario, we get flatter SMFs at the low-mass end. The effect gets more drastic when we add more massive present-day substructures since we increase the amplitude more towards the high-mass end than in the low-mass end.
VI.2.2 Increasing the Present-day Stellar-Mass Cutoff for Intact Galaxies
We can reliably select present-day intact galaxies down to $M_{\star}\gtrsim 10^{4.5}\,M_{\odot}$, but disrupted galaxies only down to $M_{\star}\gtrsim 10^{6}\,M_{\odot}$. So we test our results by also selecting intact present-day galaxies with $M_{\star}\gtrsim 10^{6}\,M_{\odot}$. When we use progenitor galaxies with $M_{\star}\gtrsim 10^{4.5}\,M_{\odot}$ at high redshifts, we find similar results for the proto-MW/LG slope in the three samples, with the least halo-to-halo scatter in case of Sample III in the all-progenitors scenario. However, in the fossil-record scenario all the three samples give a flatter slope than that of the proto-MW/LG. Upon increasing the cut-off to $M_{\star}\gtrsim 10^{5}\,M_{\odot}$ at high redshifts, the difference between the slopes in the all-progenitors scenario and the fossil-record scenario reduces, as expected. In the all-progenitors scenario the slope of Sample III is closer to the proto-MW/LG slope ($\Delta\tilde{\mu}_{\alpha}\approx 0.006$) than Sample I ($\Delta\tilde{\mu}_{\alpha}\approx 0.145$). However, in the fossil-record scenario, Sample I is more accurate ($\Delta\tilde{\mu}_{\alpha}\approx 0.007$) than Sample III ($\Delta\tilde{\mu}_{\alpha}\approx 0.150$).
@@ VI.2.3 Removing Morphologically Complex and Extremely Massive Present-day Substructures
Our present-day sample includes a few morphologically-complex, massive galaxies with $M_{\star}\gtrsim 10^{9}\,M_{\odot}$. These typically have $\sim 30-50$ progenitors and lead to more bias in the fossil-record scenario. Hence we test our results by restricting our present-day galaxy sample to those with stellar masses between $10^{6}\,M_{\odot}\lesssim M_{\star}\lesssim 10^{9}\,M_{\odot}$. The trends in this case resemble our previous results without a high-mass cut-off.
@@ VI.2.4 Lowering the High-mass Cutoff at the Present Day
We tried lowering the high-mass cutoff for present-day substructures to $M_{\star}\lesssim 10^{8}\,M_{\odot}$ while keeping the low-mass cutoff at $M_{\star}\gtrsim 10^{4.5}\,M_{\odot}$. In this case, we find that the slopes across the three samples in the all progenitors scenario, although similar, are steeper than that of the proto-MW/LG by $\approx 0.23$. In the fossil-record scenario, on the other hand, the slopes are flatter than the proto-MW/LG; with Sample III being flatter (by $\approx$0.4) than Sample II (by $\approx$0.3) or Sample I (by $\approx$0.2). Interestingly, the halo-to-halo scatter reduces only by $\sim$3% in Sample II and by $\sim$15% in Sample III. This remains similar when selecting progenitor galaxies with $M_{\star}\gtrsim 10^{5}\,M_{\odot}$ at high redshifts, although in this case all three samples produce a steeper slope than the proto-MW/LG (by  $\approx\,0.1-0.3$; Sample III is more accurate than Sample I).
@@ VI.2.5 Discussion of Sensitivity Analysis
Except when decreasing the upper mass limit, these results show that adding disrupted galaxies robustly reduces the halo-to-halo scatter, with the quantitative benefit depending on the sample selections employed. However, we can get a biased estimate of the slope in the fossil-record approach, while still reducing the halo-to-halo scatter by adding more disrupted galaxies. The bias exists even while using just the intact galaxies, with the precise bias depending, again, on sample selection. The all-progenitors scenario provides more accurate and precise recoveries of the underlying proto-MW/LG slope when disrupted galaxies are added. This motivates future efforts to combine the traditional fossil record analysis with simulated merger trees, which may allow one to properly account for the multiple high redshift progenitors of individual present-day substructures, both intact and disrupted. If, instead, only the fossil record is used, it is important to test sensitivity to the selection criteria used for both present-day and high-$z$ galaxies.
@@ VI.3 Redshift Evolution
Figure 11: Redshift evolution of the low-mass-end/faint-end slope of the SMF/UVLF. Hierarchical assembly is reflected in the increasingly steep low-mass-end/faint-end slopes with redshift for all three samples, as well as the proto-MW/LG (black) and FIREboxHR (yellow), for the differential SMF (top) and UVLF (bottom). In both panels, black, red, blue, and green points show $\tilde{\mu}_{\alpha}$ values for the proto-MW/LG, Samples I, II, and III respectively; errorbars show the $\tilde{\sigma}_{\alpha}$ values as a measure of halo-to-halo scatter. Empty markers indicate slopes estimated in the all-progenitors scenario, while filled markers denote the fossil-record case.
@@
We also estimate $\alpha$ for the SMFs/UVLFs at $z=$ 6, 8, and 9, using our fiducial choice of $M_{\star}\geq 10^{5}\,M_{\odot}$ or $M_{UV}\leq-10$ at high redshifts. We show the $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$ statistics as our central estimates for the low-mass-end/faint-end slope and the halo-to-halo scatter, respectively. The top panel of Figure 11 shows results for the low-mass-end slopes of the SMFs and the bottom panel shows the faint-end slopes for the UVLFs. The slope estimates from the all-progenitors scenario is shown by the empty markers and dashed errorbars, while the filled markers and solid errorbars denote the slope estimates from the fossil-record scenario. As a consequence of hierarchical structure formation in the cold dark matter picture with standard radiative and supernova feedback, the SMFs/UVLFs should generally get steeper towards higher redshifts. This is well-reflected by the SMFs/UVLFs from FIREboxHR (yellow), the proto-MW/LG (black), and all the three samples (I - red; II - blue; III - green).
@@
In general, Sample III consistently reproduces the most accurate and precise proto-MW/LG slope across the redshifts followed by Samples II and I, in the all-progenitors scenario. The trend is similar in the fossil-record scenario, except in a few cases. The reduction in the halo-to-halo scatter with Samples II and III is evident across the range of redshifts shown. In particular, at $z=$ 8 and 9, the large scatter in the slope from the progenitors of the intact galaxies alone is fairly well mitigated upon adding the disrupted galaxies, as shown by Sample III. We caution, however, that the $z=8$ and $z=9$ results may be subject to the limited statistical samples of progenitor galaxies captured in our simulations at these redshifts. At $z=$ 9, the faint-end slopes from our reconstructed UVLFs are shallower than the universal one. These estimates are probably poor because of the low halo abundance at these redshifts. Overall, for both the SMFs/UVLFs and for most redshifts, Sample III reproduces the proto-MW/LG slope more accurately and also reduces the halo-to-halo scatter, followed by Samples II and I. This is true for both the all-progenitors and fossil-record scenarios using our fiducial selection criteria. However, as discussed in section VI.1, the results can vary depending on our sample selection criteria, both at the present-day and at high redshifts.
VII Reconstructed Cosmic Ultraviolet Density
@@
For each of the three samples in each halo, we compute the reconstructed cosmic UV density and explore their contribution to the total cosmic UV density. As shown in Figure 12 (top panel), the proto-MW/LG (which also includes the progenitors of the present-day main host galaxy), can recover between $\sim$18% and $\sim$45% of the total universal cosmic UV density. We cannot fully recover the total universal cosmic UV density because of the difference in normalization between the UVLFs here and that of FIREboxHR, as previously discussed in section V. Hence, we mainly compare the contribution from the three samples and the proto-MW/LG. Sample III ($\sim$8–18%) is the most complete, followed by Samples II ($\sim$4–16%), and sample I ($\sim$1–8%).
We also consider the relative contribution of different progenitor galaxies to the total reconstructed cosmic UV density within each sample at $z=7$ (Figure 12, bottom panel). The median UV luminosity density reconstructed from Sample I (red) under-predicts the contributions from galaxies brighter than $M_{\mathrm{UV}}\sim-15$, which is the current JWST detection limit for unlensed galaxies  [^5] Measurements behind foreground lensing clusters can probe fainter galaxies down to $M_{\mathrm{UV}}\gtrsim-13$ (Chemerynska et al., 2025), but they come with systematic uncertainties.. Furthermore, the large scatter in the faint-end slope of the reconstructed UVLF from Sample I also leads to sizable uncertainties in the inferred UV luminosity density and the relative importance of different galaxies in reionization. Quantitatively, as inferred from Sample I, the luminosity above which galaxies produce $\sim 50\%$ of the total UV density is somewhere in the range $16\lesssim M_{\mathrm{UV}}\lesssim-13.5$. Adding the progenitors of the disrupted galaxies in Samples II and III (blue and green shaded regions, respectively) reduces the scatter. Sample II predicts minimum UV magnitudes in the range $16.5\lesssim M_{\mathrm{UV}}\lesssim-14.5$, while Sample III predicts a range of $16.5\lesssim M_{\mathrm{UV}}\lesssim-15$. Furthermore, the median curves (solid curves) are more representative of the proto-MW/LG curve (black) when the disrupted galaxies are included. The median curve for Sample III (green) shows slightly larger contributions from brighter galaxies than in the proto-MW/LG, because it predicts a shallower faint-end slope than that of the proto-MW/LG in the fossil-record scenario (see Figure 10).
@@ VIII Detectability of the Present-day Substructures
To complete our analysis, we construct UVLFs by imposing the detectability constraints obtained from the internal view scenario using LSST. Figure 4—first shown in Section V.1—gives the differential UVLFs at $z=7$ reconstructed from different classification groups (solid curves) in comparison to reconstructions from their detectable counterparts alone (dashed curves). The reconstructed UVLFs appear promising, even after imposing the detectability constraints. Although the normalization of the satellite and field galaxy progenitor UVLFs is reduced, their shapes are relatively well-preserved. The progenitor UVLFs reconstructed from the streams remain almost unchanged even with just $\sim$50 % detectability, since the detectable streams are mostly of higher mass and each massive stream typically has numerous reionization-era progenitors (see Section IV.2), spanning much of the high-$z$ distribution in progenitor luminosity. We show the results in the all-progenitors scenario by utilizing the full merger histories of the present-day substructures, and not in the fossil-record scenario. However, the difference between the shape of the UVLFs with and without detectability constraints does not change significantly in the fossil-record scenario.
@@ IX Summary
In this work, we test whether including present-day disrupted galaxies like stellar streams and phase-mixed galaxies in the near-far technique improves the inference of the low-mass-end/faint-end slope of the SMF/UVLF during the EoR. Using the FIRE-2 cosmological zoom-in baryonic simulations of MW-like (isolated) and LG-like (paired) halos (Section II), we classify present-day substructures into 4 categories: intact satellite galaxies, field galaxies, stellar streams, and phase-mixed galaxies (Section III). The first two categories constitute our intact galaxies sample (Sample I); the intact galaxies and streams comprise our Sample II; adding the phase-mixed galaxies makes our Sample III. Using particle tracking, we assign each star particle to its progenitor galaxy at high redshifts. We then assess the contribution of different progenitor populations to the total UV luminosity density (Sections IVVI). Finally, we estimate the proportion of substructures that will be detected by LSST (for the Milky Way) or Roman (for external galaxies) and the effect of these detection limits on the reconstructions (Section VIII). Our results are summarized below.
Our results are consistent in general with Gandhi et al. (2024), with a few differences. Gandhi et al. (2024) estimated the low-mass-end slopes of the reconstructed fossil record SMFs using just intact low-mass galaxies and found close agreement with the proto-MW/LG slopes. We reach similar conclusions but also show that the halo-to-halo scatter associated with the slope is large when using just intact galaxies. Their slope estimates are not exactly the same as our estimates for Sample I, since some of their present-day low-mass galaxies are part of our streams and hence do not show up in our Sample I. This also leads to differences in the shape and normalization of our respective intact low-mass galaxy fossil-record SMFs, with a stronger mass dependent bias in our case. Additionally, our methods for computing the slopes are different. Furthermore, Gandhi et al. (in prep.), use an alternative set of zoom-in high-redshift FIRE-2 simulations as a benchmark for the universal SMFs/UVLFs, instead of FIREboxHR. Both sets of simulations show good agreement with each other, as well as with observed high-$z$ SMFs/UVLFs and cosmic UV luminosity density, but they may not match exactly.
@@
Inclusion of tidally disrupted galaxies leads to more accurate high-$z$ reconstructed SMFs/UVLFs. We analyze the SMFs/UVLFs under two scenarios: the all-progenitors scenario, in which we account for the stellar mass contribution from each progenitor of a present-day substructure separately; and the fossil-record scenario, where we assume that the stellar mass from all progenitors of each substructure is contained within a single galaxy, as would be inferred observationally in the absence of merger history information. In both of these scenarios, our Sample III recovers most of the amplitude ($\sim$75%) of the proto-MW/LG SMF/UVLF, followed by Samples II ($\sim$50%) and I ($\sim$25%), respectively. This represents a significant improvement, reducing the normalization rescaling from a factor of $\sim$4 in Sample I to just $\sim$1.5 in Sample III, after including the disrupted galaxies. However, in the fossil-record scenario, the SMFs/UVLFs become shallower in the low-mass-end/faint-end region, with the effect being large when more disrupted galaxies are added. Nevertheless, in most instances, the SMFs and UVLFs in the fossil-record scenario trace the shape of the proto-MW/LG SMFs/UVLFs between $10^{5}\,M_{\odot}\lesssim M_{\star}\lesssim 10^{6.5}\,M_{\odot}$ and $14\lesssim M_{\mathrm{UV}}\lesssim-10$, respectively. The fossil-record results differ from the all-progenitors case because the latter approach combines multiple low-mass/faint progenitors into a single more massive system. This difference is compounded in the case of disrupted galaxies which are generally more massive and have multiple reionization-progenitors. We also compare our results with the differential SMFs/UVLFs of the FIREboxHR simulation, as a benchmark for the universal SMF/UVLF, and find close agreement in shape. Although the proto-MW/LG slopes are in agreement with FIREboxHR within our estimated uncertainties, they are in general flatter by $\sim$0.07. This could be due the difference in the mean matter density of the region around the MW/LG and the mean matter density of the universe, but a detailed exploration is needed, which is beyond the scope of this paper.
@@
The low-mass-end/faint-end slope is more robustly constrained when including disrupted galaxies. We estimate the low-mass-end/faint-end slope of the reconstructed SMFs/UVLFs using a Bayesian approach, to test how accurately and precisely we can recover the proto-MW/LG and the universal SMF/UVLF slopes. In the all-progenitors scenario, adding disrupted galaxies helps reproduce the low-mass/faint-end slope of the proto-MW/LG SMF/UVLF more accurately and more precisely, and these are also representative of the global-average low-mass/faint-end slopes. Sample III reproduces the proto-MW/LG slope most accurately with an offset of $\approx$0.04, followed by Samples II ($\approx$0.12) and I ($\approx$0.26). Furthermore, when we use only the intact galaxies, there is a large halo-to-halo scatter due to the varying accretion histories across different halos, even though the median values agree well with that from the proto-MW/LG. Sample III reduces the scatter of Sample I by $\sim$40%, while Sample II reduces it by $\sim$20%. Including the disrupted galaxies improves our inferences of the median slopes and also reduces the halo-to-halo scatter. In the fossil-record scenario, we find similar results when we select high-$z$ progenitors with $M_{\star}\geq 10^{5}\,M_{\odot}$ and $M_{\mathrm{UV}}\leq-10$. However, even though the halo-to-halo scatter decreases, we can get a biased estimate of the slope from the fossil record depending on our choice of both the present-day and high-$z$ galaxy samples. The estimated slopes vary depending on the stellar mass range of the galaxies selected, both at the present-day and during the EoR, even in the case of using only intact galaxies. In the future, it may be possible to augment the observed SFHs with average merger histories from cosmological simulations, providing yet more accurate and precise estimates of the low-mass-end/faint-end slope. Nevertheless, including the streams and phase-mixed galaxies improves the statistical constraints on the proto-MW/LG SMFs/UVLFs, reducing the halo-to-halo scatter ($\sim$0.13 in case of proto-MW/LG) from $\sim$0.24 (Sample I) to $\sim$0.20 (Sample II) and $\sim$0.15 (Sample III). A similar trend holds when we separate into MW-like and LG-like environment. This improvement is therefore crucial, since we have only one MW and LG with which to apply the near-far technique.